OpenStudy (anonymous):

What is the factored form of the expression? w^2+18w+77 A. (w-7)(w+11) B. (w-7)(w-11) C. (w+7)(w+11) D. (w+1)(w+77)

OpenStudy (anonymous):

D :)

OpenStudy (anonymous):

C

OpenStudy (anonymous):

What times what is 77 but when added = 18?

OpenStudy (anonymous):

C*

OpenStudy (anonymous):

sorry its C :)

OpenStudy (anonymous):

... don't just give the answer.... serious, how does that help anyone?

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (anonymous):

@agentc0re i m not going to answer that :P

OpenStudy (anonymous):

LOL @ ^

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

if you have anymore just post it and make sure to tag me in it :)

OpenStudy (anonymous):

how do you tag

OpenStudy (anonymous):

@ + yummydum = @yummydum :)

OpenStudy (anonymous):

alright :) bout to post another one

OpenStudy (anonymous):

tagggg :D

OpenStudy (anonymous):

well you're probably going to get screwed on an exam if you don't know how to do it then. Math is extremely important, even if you don't realize it right now. :/

OpenStudy (anonymous):

okay here....just so agentc0re can go to sleep at night -_- to factor an expression you need to look at th coefficient of the 3rd term....which two of its factors add up to the coefficient of the second term example: x^2+5x+6 2 * 3 = 6 2 + 3 = 5 now that you've found those numbers...make two binomials as such: (x+__)(x+__) then you put in the 2 numbers you just found (x+2)(x+3) x^2+5x+6 = (x+2)(x+3)